Find the maximum value of the function f(x) = x + x2
1
Expert's answer
2019-08-05T13:13:18-0400

Answer: there are no local maximum points.


"supf(x)=+\infty"

Explanation. The function f(x) is unbounded above, since

"\lim_{x\to \infty }[(x^{2}+x+\frac{1}{4})-\frac{1}{4}]=\lim_{x\to\infty}[(x+\frac{1}{2})^{2}-\frac{1}{4}]=+\infty"

The only stationary point

"x=-\frac{1}{2} , 
(f'(-\frac{1}{2})=0)"

is the local minimum point, as can be seen from the inequality:

"f(x)=x+x^{2}=(x+\frac{1}{2})^{2}-\frac{1}{4}\geqslant -\frac{1}{4}=f(-\frac{1}{2})"

There are no maximum points